A rifle is found to give the bullet a muzzle velocity of 485 m/s the barrel is 80cm long
a) how much time is the bullet moving in the barrel?
b) what is the average acceleration of the bullet?
1
Expert's answer
2019-10-14T09:13:10-0400

The movement of the ball is evenly accelerated.

We can write


"2as=v^2-{v_0}^2 (1)"


In our case, v0=0m/s, v=485 m/s, s=0.8 m

Using (1) we get

"a=\frac{ v^2}{2s} (2)"


We have

a=147015 m/s2


The acceleration is given by formula

"a=\frac{v-v_0}{t} (3)"


Using (3) we get

"t=\frac{v-v_0}{a} (4)"


Answer

a)     0.0033 s

b)     147015 m/s2